OpenStudy (anonymous):

a rectangle of perimeter 100 units has the dimensions shown. Its area is given by the function A = w(50 - w). What is the GREATEST area such a rectangle can have?

OpenStudy (jonnyvonny):

We know the formula for the perimeter of a rectangle: P=2l+2w. We're given the width, so lets get width alone:\[200=2l+2w \rightarrow 100=l+w \rightarrow 100-l=w\] substitute this equation with the one given: A=w(50-w): \[100-l=w \rightarrow A=w(50-w)\rightarrow A=(100-l)(50-(100-l))\]\[A=(100-l)(50-100+l)\rightarrow A=(100-l)(-50+l)\]\[A=(100-l)(-50+l)\rightarrow A=-5000+100l+50l+l^2\]\[A=-5000+150l+l^2\] From there, either factor or get 2 binomials, then cross out that gives a negative x-value. ie. x=-200 or something.

OpenStudy (anonymous):

Do you still need help, if not you should close the question.

OpenStudy (jonnyvonny):

^